Integrate sin2x (plus) cos2x ,limit x=0 to x= pi
Answer by Alan3354(69443)

If you mean
sin(2x) + cos(2x)
-----
INT = -cos(2x)/2 + sin(2x)/2
INT(pi) = -1/2 + 0
INT(0) = -1/2 + 0
-----------------
--> 0